OpenStudy (anonymous):

Solve for x: –4(3x – 2) = 6x + 2

OpenStudy (anonymous):

\[-4\times3x=-12x, -4\times-2=8, -12x+8=6x+2\]

OpenStudy (anonymous):

Can you solve from there, or do you want me to help?

OpenStudy (anonymous):

@Its_James0808

OpenStudy (anonymous):

Whats the answere @rileychampion365

OpenStudy (anonymous):

Lol, okay then. \[-12x+8=6x+2, Add 12x \to both sides, 8=12x+2, Subtract 2 from both sides, 6=12x, Simplify, 12\div6. X=2\]

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

−12x+8=6x+2,Add12x→bothsides, 8=12x+2,Subtract2frombothsides,6=12x, Simplify,12÷6.X=2